OpenStudy (anonymous):

The curve y=x^2+ax+b and y=cx-x^2 have common tangent line at (1,0) find a ,b, c?

OpenStudy (anonymous):

i have try to solve but at end i am confused please solve it. i have find tangents of both curves and equate them,then find the equation of tangent which must be equall but its not coming equal diffrence in x^2 value sign

OpenStudy (anonymous):

dont equate the equation of tangents.. if two equations are identical then ratio of coefficients are equal

OpenStudy (anonymous):

suppose a1x+b1y+c1=0 and a2x+b2y+c2=0 are identical equations then a1/a2=b1/b2=c1/c2

OpenStudy (shubhamsrg):

i assume it means both curves pass through (1,0) and also just touch each other at that point so you'll have 1+a+b = 0 and c-1 = 0 tangent at x=1 for second eqn = c- 2 and for 1st one is 2+a => c-2 = 2+a you may now easily solve for a,b,c.. i may have mistreated the question though,please correct me of am wrong somewhere..